### 25 Mar 2026 v1

# A New Method to Simulate Dark Matter–Baryon Interactions and Application to an Isolated Disk Galaxy

Abstract
We report on a new method for incorporating interactions between dark matter (DM) and baryons in cosmological simulations, capable of handling the challenging regime in which the dark matter particle mass is comparable to or lighter than the baryon mass. The method hybridizes two distinct approaches, as gas particles receive momentum and energy transfer according to a mean-field calculation while DM particles undergo Monte Carlo scatterings, which are derived from the Boltzmann equation and proved to be statistically equivalent. We present an open-source implementation of this method in the simulation code GIZMO. As a first application, we investigate the effects of DM–baryon interactions on an isolated Milky Way-like disk galaxy for dark matter having twice the proton mass, which roughly maximizes the average energy transfer per collision. For cross sections of order 1 barn ($10^{-24}\,\mathrm{cm}^{2}$), these interactions cause strong changes to the mass distribution in the center of the galaxy in less than 1 Gyr, even when bar formation is suppressed by hand. For cross sections typical of hadronic interactions, $\lesssim 30\,{\rm mb}$, high-fidelity galaxy formation simulations will be needed to assess the effects on observable features of galaxies.
Dark matter, Hydrodynamical simulations, N-body simulations, Galaxy dynamics, Particle astrophysics
I Introduction
[…]
The most commonly known particle physics model featuring non-gravitational interactions between DM and the Standard Model is the Weakly Interacting Massive Particle (WIMP). WIMP dark matter is a natural prediction of supersymmetry, but the parameter space for WIMP dark matter is now severely constrained by accelerator limits and direct detection experiments such as Aprile et al. (2023); Aalbers et al. (2025). Moreover, due to its heavy mass and small cross section with baryons, WIMP dark matter is effectively non-interacting except gravitationally.
Here, we are interested in dark matter which has observationally impactful interactions with baryons. Direct detection of DM which is light and has moderate or strong interactions with baryons is challenging. Unless it is weakly interacting, a DM particle loses most of its energy interacting in the material above the detector, leaving too little energy to be detected. Figure 8 of (Albakry and others, 2023) summarizes exclusion regions for cross sections below 1 mb ($10^{-27}\,{\rm cm}^{2}$). Xu and Farrar (2023) reviews the cross section constraints on DM in the GeV mass regime, finding that the direct detection constraints are non-existent or tentative for dark matter masses $\lesssim 3$ GeV and dark matter–baryon interaction cross sections in the $10^{-28}\text{--}10^{-25}\,{\rm cm}^{2}$ range. A proposed limit in this range, based on detection in a deep underground experiment of light DM up-scattered by cosmic rays, will potentially become constraining after non-resonant energy losses in propagation through the Earth are accounted for (Alvey et al., 2023). Limits are also obtained from astrophysical systems such as gas-rich dwarf galaxies, cf., Wadekar and Farrar (2021), clusters, and many others.
A complementary approach to constraining dark matter–baryon (DM–b) interactions is through their astrophysical impacts on the structure of galaxies (e.g., Dvorkin et al., 2014; Buen-Abad et al., 2022). The effects on the structure of galaxies come about in two primary ways: by introducing viscosity before recombination, suppressing small scale power, and by changing the formation and evolution of galaxies. The pre-recombination imprints have been used to place limits on interaction parameters by comparing against measurements of the CMB (Gluscevic and Boddy, 2018; Boddy and Gluscevic, 2018; Boddy et al., 2018; Ali-Haïmoud et al., 2024), the Lyman-$\alpha$ forest (Dvorkin et al., 2014; Xu et al., 2018), large-scale structure (He et al., 2023, 2025), and the abundance of Milky Way satellites (Nadler et al., 2021; Maamari et al., 2021; Buen-Abad et al., 2022).
[…]
The purpose of our work is to develop tools to enable an accurate treatment of the effects of dark matter–baryon interactions during galaxy formation and evolution. We have developed a new method for incorporating DM–b interactions into hydrodynamic cosmological simulations, which we have implemented as a new module in the code GIZMO (Hopkins, 2015). The present implementation treats elastic scatterings between non-relativistic dark matter and baryon particles with a phenomenological, isotropic power-law cross section $\sigma(v)=\sigma_{n}\,v^{n}$, with $v\ll c$ the relative velocity, however the method applies for general cross section dependencies on $v$ and scattering angle. The method hybridizes two statistically equivalent but qualitatively different approaches: the dark matter is treated in a Monte Carlo fashion, where simulation particles probabilistically undergo scatterings, and the baryons are treated with a mean field approach, where the statistical expectations for the transfer rates of bulk velocity and thermal energy are applied. These are both derived directly from the Boltzmann equation, ensuring that energy and momentum transfer between the sectors are equal and opposite on average.
During the preparation of this article, we learned of a parallel effort to implement dark matter–baryon interactions in $N$-body simulation from Fischer et al. (2025). There, the simulation dark matter particles scatter off of “virtual” baryon particles, with the resulting energy and momentum change of the virtual particle distributed back onto the actual simulation baryon macroparticles. While we defer a more detailed comparison with their method to the main text, we note that there are two important differences between our methods. The Fischer et al. (2025) approach conserves energy and momentum exactly per scattering, while we give this up and conserve energy and momentum only on average. However, their virtual particle scheme requires that the masses of the simulation macroparticles are in a similar ratio as the masses of the elementary dark matter to protons. In high-quality simulations of galaxies, much less resolution is needed for the DM particles than for baryons without DM–b interactions, and DM macroparticles masses can be orders of magnitude heavier than those for baryons. Thus to apply the F25 method with comparable mass DM and baryon macroparticles may be computationally infeasible for $m_{\chi}$ comparable to or less than the mass of the proton. By contrast, our approach—by treating the baryons separately from the probabilistic DM–b scattering—sidesteps this issue entirely, enabling us to press into light dark matter masses without huge computational overhead.
Our method thus provides for the first time a practical tool designed for investigating dark matter–baryon interactions with dark matter mass close to that of hydrogen and helium, where energy and momentum transfer in individual collisions is maximally efficient. A concrete example of such a model is Sexaquark Dark Matter, in which the dark matter particle is a six-quark state (uuddssuuddss) whose binding is sufficiently deep that it is effectively stable; see Farrar (2022) and references therein for a review of the constraints from particle physics, astrophysics, and cosmology.
As an initial application of our code, we simulate the isolated Milky Way-like disk galaxy of the AGORA project (Kim et al., 2016) with a dark matter particle having twice the proton mass ($2\,m_{p}$) and velocity-independent cross section. We find that sufficiently strong dark matter–baryon interactions induce significant changes to the morphology of the center of the galaxy. In particular, such interactions drive gas and dark matter inflows toward the center. For potentially realistic DM–baryon cross sections, below tens of millibarns ($10\ {\rm mb}=10^{-26}\ {\rm cm}^{2}$), assessing the impact of dark matter–baryon interactions will require more detailed treatment of star formation and feedback, and of the hierarchical formation of galaxies, than we undertake in the present work.
The article is structured as follows. Section II gives the relevant scattering theory and the algorithms used to implement our approach in code. Section III briefly discusses the tests we perform to validate our implementation. Section IV presents the key results of our simulation of the AGORA disk galaxy. Section V concludes. Appendices provide derivations and details about our test suite.
II Method
II.1 Boltzmann equation
We consider elastic $2\leftrightarrow 2$ collisions between non-relativistic dark matter particles (of mass $m_{\chi}$) and baryons (of mass $m_{B}$) with differential cross section ${\rm d}\sigma/{\rm d}\Omega(v_{\rm rel},\theta)$. In any individual scatter with initial velocities $\bm{v}_{\chi},\bm{v}_{B}$, the outgoing velocities are given by
[…]
where $\bm{v}_{\chi B}\equiv\bm{v}_{\chi}-\bm{v}_{B}$, $v_{\chi B}=|\bm{v}_{\chi B}|$, $\hat{\bm{v}}_{\chi B}=\bm{v}_{\chi B}/|\bm{v}_{\chi B}|$, and $\hat{\bm{v}}_{\chi B}^{\prime}$ is a new direction. The scattering angle $\theta$ describes the angle between the two unit vectors $\hat{\bm{v}}_{\chi B}$ and $\hat{\bm{v}}_{\chi B}^{\prime}$: $\cos\theta=\hat{\bm{v}}_{\chi B}\cdot\hat{\bm{v}}_{\chi B}^{\prime}$. We introduce $\bm{S}_{\chi}$ and $\bm{S}_{B}$, which may be thought of as functions of $\bm{v}_{\chi},\bm{v}_{B},\Omega$, as a short notation representing this transformation.
In the continuum limit, dark matter and baryons are each described by a distribution function [^1] We use the convention $\int{\rm d}^{3}x\,{\rm d}^{3}v\,f(\bm{x},\bm{v})=\mathcal{N}$, the total number of particles. The number density is then given by $n(\bm{x})=\int{\rm d}^{3}v\,f(\bm{x},\bm{v})$. (df) $f(\bm{x},\bm{v})$ whose evolution is governed by the Boltzmann equation:
$\frac{{\rm d}f_{\chi}}{{\rm d}t}(\bm{v}_{\chi})$ $=\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,$ (3)
$\qquad\times\big[f_{\chi}(\bm{S}_{\chi})\,f_{B}(\bm{S}_{B})-f_{\chi}(\bm{v}_{\chi})\,f_{B}(\bm{v}_{B})\big],$
where the dependence on $\bm{x}$ is implicit. The right-hand-side of this equation is known as the collision integral.
Our goal is to implement the effects of these collisions in a Lagrangian-picture hydrodynamic cosmological simulation. Such simulations are built on three approximations, which we will use to derive our method.
The first approximation made by all $N$-body simulations is that the dark matter is a “softened $N$-body component”; namely, that a number of Monte Carlo samples are drawn from some true initial df $f_{\chi}$ and spatially convolved with a smoothing kernel W(r)W(r), [^2] We assume throughout that any valid smoothing kernel $W(r,h)$ integrates to 1 and is zero (or else negligible) for $r>h$. producing the following approximate df
$\hat{f}_{\chi}(\bm{x},\bm{v})=\sum_{i}\frac{M_{i}}{m_{\chi}}\,W(|\bm{x}-\bm{x}_{i}|,h_{i})\,\delta(\bm{v}-\bm{v}_{i}).$ (4)
This can be understood as representing the dark matter by $N_{\chi}$ “macroparticles”, each having a mass $M_{i}$, position $\bm{x}_{i}$, smoothing length $h_{i}$, and velocity $\bm{v}_{i}$.
The second approximation is that the gas is represented by a thermalized softened $N$-body component. By this, we mean that the gas is decomposed into macroparticles like dark matter, but each macroparticle has a bulk velocity and internal energy describing the local velocity distribution, which is taken to be Maxwell-Boltzmann at all times. Therefore, the gas is represented by the approximate df
$\hat{f}_{B}(\bm{x},\bm{v})=\sum_{j}\frac{M_{j}}{\mu_{j}}\,W(|\bm{x}-\bm{x}_{j}|,h_{j})\,\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j}),$ (5)
where $\mathcal{G}$ is the probability density function of a three-dimensional isotropic Gaussian—i.e., the Maxwell-Boltzmann distribution. In simulation, then, the gas is represented by $N_{B}$ macroparticles, each having a mass $M_{j}$, position $\bm{x}_{j}$, smoothing length $h_{j}$, bulk velocity $\bm{V}_{j}$, temperature $T_{j}$, and mean molecular weight $\mu_{j}$.
We note that the form of the baryon df given in Equation 5 is valid regardless of the choice of hydrodynamic solver. By computing the moments of the Boltzmann equation for exactly this df, one can derive the fluid equations for standard smoothed particle hydrodynamics (e.g., Monaghan, 1992). More modern solvers, such as the moving Voronoi mesh of AREPO (Springel, 2010) or the meshless methods of GIZMO (Hopkins, 2015), generally assume sophisticated geometries for discretizing the gas. However, in these simulations the resulting particle distribution is then taken to have the form of Equation 5 for other applications, such as solving gravity, and so we take this form for solving interactions with dark matter.
The macroparticle dfs for dark matter and gas have a similar form. For brevity, we introduce the following notation:
$\hat{f}(\bm{x},\bm{v})=\sum_{i}n_{i}(\bm{x})\,\nu_{i}(\bm{v}),$ (6)
where $n_{i}$ is a number density and $\nu_{i}$ is a probability density function for the velocities, equal to either a Dirac delta (for dark matter) or a Maxwell-Boltzmann distribution (for gas).
The third assumption underpinning these simulations is that the time-evolution of the true dfs $f_{\chi},f_{B}$ is well-approximated even when enforcing that the form of the macroparticle dfs $\hat{f}_{\chi},\hat{f}_{B}$ is preserved. In other words, after a timestep, the forms of Equations 4 and 5 can be imposed on the dark matter and gas, albeit with updated values of the parameters. We make this point explicit because the full evolution of the df according to the Boltzmann equation would distort the functional form of the discretization—for example, the softened macroparticles would experience tidal forces and become stretched and sheared—but this is regularly ignored. For our purposes, this assumption means we can impose that evolution according to the collision integral does not change the form of the discretization: after an update, the dark matter will have the form Equation 4 and the baryons will have the form Equation 5. Because the collision integral is local, it will also not change the masses or positions of any particles. Thus, we need only consider the update to the individual particle velocity distributions, and so the Boltzmann equation is
$\sum_{i}n_{i}(\bm{x})\,\frac{{\rm d}\nu_{i}}{{\rm d}t}\bigg|_{\bm{v}_{\chi}}$ (7)
$\qquad=\sum_{i,j}n_{i}(\bm{x})\,n_{j}(\bm{x})\,\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,$
$\qquad\qquad\times\big[\nu_{i}(\bm{S}_{\chi})\,\nu_{j}(\bm{S}_{B})-\nu_{i}(\bm{v}_{\chi})\,\nu_{j}(\bm{v}_{B})\big],$
and similarly for the baryons $\hat{f}_{B}$.
The form of this equation suggests a matching procedure: term $i$ on the left-hand-side can be set equal to term $i$ on the right-hand-side. By doing this matching, we are making an additional assumption: that the phase space densities of macroparticles of the same type do not mix, so each macroparticle evolves as if it is the only particle of its type at its location. This assumption is reasonable as long as no particles have identical positions $\bm{x}$, which is already required for the validity of the $N$-body approach. Hence, individual macroparticles evolve according to independent collision integrals:
$\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}(\bm{v}_{\chi})$ $=\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,$ (8)
$\qquad\times\big[f_{\chi}^{i}(\bm{S}_{\chi})\,\hat{f}_{B}(\bm{S}_{B})-f_{\chi}^{i}(\bm{v}_{\chi})\,\hat{f}_{B}(\bm{v}_{B})\big],$
with $\bm{x}$-dependence implicit and defining $f_{\chi}^{i}(\bm{x},\bm{v})\equiv n_{i}(\bm{x})\,\nu_{i}(\bm{v})$. This equation applies analogously to the baryons.
II.2 Dark matter evolution
In order to determine how to update the velocities of the dark matter particles, consider a finite timestep $\Delta t$. To first order in the timestep, the updated df is
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t$ (9)
where we require $f_{\chi}^{i}\big|_{t+\Delta t}$ to have the form $n_{i}(\bm{x})\,\nu_{i}^{\prime}(\bm{v})$, for $\nu_{i}^{\prime}$ the new velocity distribution of the particle.
The derivation proceeds from here in detail in Appendix A. Briefly, the steps we take are as follows. We substitute into Equation 9 the baryon macroparticle df $\hat{f}_{B}$. This gives a sum over gas particles $j$, each with a different integral over ${\rm d}^{3}\bm{v}_{B}\,{\rm d}\Omega$. We simulate this integral by drawing a single Monte Carlo sample independently for every $j$ (denoting the sampled baryon velocity $\bm{v}^{B}_{ij}$ and scattering angles $\Omega_{ij}$). Then, in order to isolate the velocity distribution on both sides, we integrate both sides over all space. The result is
$\nu_{i}^{\prime}(\bm{v})=\Big(1-\sum_{j}P_{ij}\Big)\,\delta(\bm{v}-\bm{v}_{i})+\sum_{j}P_{ij}\,\delta\big(\bm{v}-\bm{S}_{\chi}(\bm{v}_{i},\bm{v}_{ij}^{B},\Omega_{ij})\big),\quad\text{where}$ (10)
$P_{ij}\equiv\frac{M_{j}}{\mu_{j}}\,g_{ij}\,\sigma(v_{\rm rel})\,v_{\rm rel}\,\Delta t,\quad g_{ij}\equiv\int{\rm d}^{3}x\,W(|\bm{x}-\bm{x}_{i}|,h_{i})\,W(|\bm{x}-\bm{x}_{j}|,h_{j}),\quad\sigma(v)\equiv\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\bigg|_{v},$ (11)
and $v_{\rm rel}\equiv|\bm{v}_{i}-\bm{v}_{ij}^{B}|$. Note the appearance of the total cross section $\sigma(v)$. We have introduced two intermediate variables to clarify the meaning of this result. One is $g_{ij}$, introduced by Rocha et al. (2013) for self-interacting dark matter simulations and dubbed the “geometric factor” or the “overlap integral”. It has dimensions of inverse volume and gives the effective number density of particle $j$ acting on particle $i$. The other is $P_{ij}$ which is dimensionless and, as long as $P_{ij}\ll 1$, can be understood as a probability. Indeed, $(M_{j}/\mu_{j})\,g_{ij}$ is the effective number density of microscopic baryons active in this interaction, so $P_{ij}$ has the form $n\,\sigma\,v\,\Delta t$, exactly as one would expect.
The new velocity distribution $\nu_{i}^{\prime}$ is a discrete probability distribution: it gives the post-scattering velocity $\bm{S}_{\chi}(\bm{v}_{i},\bm{v}^{B}_{ij},\Omega_{ij})$ with probability $P_{ij}$, or the original velocity $\bm{v}_{i}$ with probability $1-\sum_{j}P_{ij}$. To pick a new velocity for macroparticle $i$, then, we can draw a sample from this discrete distribution. Note that the time-evolution of the dark matter macroparticles is given by probabilistically scattering them off of samples from the baryon velocity distribution in a manner which agrees exactly with the microscopic theory.
II.3 Baryon evolution
We next consider the updates to apply to the baryons. As noted above, we require that the updated baryon velocity distribution remain Maxwell-Boltzmann. This suggests an alternative approach in which we directly compute the rates-of-change of the parameters of the Maxwell-Boltzmann distribution: the bulk velocity and the temperature. These are
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,\frac{{\rm d}f_{B}^{j}}{{\rm d}t}\bigg|_{\bm{x},\bm{v}_{B}}\,\bm{v}_{B},$ (12)
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,\frac{{\rm d}f_{B}^{j}}{{\rm d}t}\bigg|_{\bm{x},\bm{v}_{B}}\,(\bm{v}_{B}-\bm{V}_{j})^{2}.$ (13)
These can be evaluated using techniques very similar to those described by Ali-Haïmoud (2019), augmented by an integral over space as in the dark matter case. Details are provided in Appendix B. The result is
$\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $\approx\sum_{i}\frac{M_{i}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\,(\bm{v}_{i}-\bm{V}_{j}),$ (14)
$\frac{3}{2\,\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $\approx\sum_{i}\frac{M_{i}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\bigg[(\bm{v}_{i}-\bm{V}_{j})^{2}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg],$ (15)
where $g_{ij}$ is the overlap integral as before and $\mathcal{A}$ and $\mathcal{B}$ are defined
$\mathcal{A}(\bm{w};\varsigma^{2})$ $\equiv\int{\rm d}^{3}u\,\mathcal{G}(\bm{w}-\bm{u},\varsigma^{2})\,u\,\frac{\bm{u}\cdot\bm{w}}{w^{2}}\,\overline{\sigma}(u),$ (16)
$\mathcal{B}(\bm{w};\varsigma^{2})$ $\equiv\int{\rm d}^{3}u\,\mathcal{G}(\bm{w}-\bm{u},\varsigma^{2})\,u^{3}\,\overline{\sigma}(u),$ (17)
$\overline{\sigma}(v)$ $=\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\bigg|_{v}\,(1-\cos\theta).$ (18)
Notice the appearance of the momentum transfer cross section, $\overline{\sigma}$. If the momentum transfer cross section has a power law dependence on $v$, then these integrals $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions, given in Appendix B. In simulation, these rates-of-change can simply be treated as new sources of acceleration and heating, added onto the total acceleration and heating rates which are already accumulated for each particle and integrated by the preferred scheme of the simulation.
These rates-of-change are statistically equal-and-opposite to the Monte Carlo sampling procedure we have outlined for the dark matter. Momentum and energy are conserved in expectation value, which we prove in Appendix C.
We note that the dark matter and baryon calculations end up using different integrals over the differential cross section: the dark matter scattering probability depends on the total cross section $\sigma$, while the gas momentum and heat exchange rates depend on the momentum transfer cross section $\overline{\sigma}$. In the self-interacting dark matter literature, it is often assumed that one can approximate a non-isotropic cross section in the scattering probability by using isotropic scatterings with the momentum transfer cross section (Tulin and Yu, 2018). For dark matter–baryon interactions, such an assumption must be carefully considered, as it could break the equivalence between the two calculations.
II.4 GIZMO implementation
We implement this method in GIZMO (Hopkins, 2015), which is based in part on the algorithms of GADGET-2 (Springel, 2005). We choose to specialize to an isotropic power-law cross section $\sigma(v)=\overline{\sigma}(v)=\sigma_{0}\,(v/c)^{n}$. We note that other cross sections, including non-isotropic ones, can be implemented in this framework with relative ease, so long as the integrals defining $\mathcal{A}$ and $\mathcal{B}$ can be computed efficiently.
The core of the implementation is a nearest-neighbors calculation, where dark matter particles seek gas particles and vice versa. Each particle searches in a radius equal to twice its softening length. With this search radius, each pair of particles which overlaps is guaranteed to be found by the neighbor with larger softening length, but may or may not be found by the smaller neighbor. As such, we must search in both directions (dark matter $\rightarrow$ gas and vice versa); to ensure that each pair is processed only once, we skip a pair whenever the searching particle has the smaller softening length.
To process pair $(i,j)$ (taking $i$ to be dark matter, $j$ gas), we first compute the overlap integral $g_{ij}$. We then compute the contribution of particle $i$ to the acceleration ${\rm d}\bm{V}_{j}/{\rm d}t$ and heating rate ${\rm d}T_{j}/{\rm d}t$ of particle $j$, which are accumulated over all pairs that include $j$. Next, we draw a random sample from the gas velocity distribution $\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j})$ and use it to compute the scattering probability $P_{ij}$. If the computed scattering probability $P_{ij}$ is too large, we reduce the timestep size for the dark matter particle for the next timestep. The specific criterion we use is
$P_{ij}<\min\bigg(0.1,10^{-4}\,\bigg[\frac{m_{\chi}+\mu_{j}}{\mu_{j}}\bigg]^{2}\bigg),$ (19)
which requires smaller timesteps if $m_{\chi}\lesssim\mu_{j}$, because in that case scatters are expected to impart a large change on the dark matter velocity. We draw a random number $r$ uniformly between 0 and 1 to decide whether to scatter the dark matter particle. If we choose to scatter (if $r<P_{ij}$), we choose a new direction $\hat{\bm{v}}_{\chi B}^{\prime}$ isotropically, then compute $\bm{v}_{\chi}^{\prime}$ from Equation 1 and apply it to particle $i$.
An important detail in the implementation is the computation of the confluent hypergeometric function $_{1}F_{1}(a,b;x)$. As given in Appendix B, the integrals $\mathcal{A}$ and $\mathcal{B}$ reduce to simple forms involving $_{1}F_{1}$ for our chosen cross section. We exclusively use this function in the regime of small $a$, small $b$, and negative $x$. However, the GSL implementation of this function is subject to overflow at modestly large values of x-x. We can improve the performance for this specific regime by making use of Kummer’s transformation and the asymptotic expansion for large values of $x$ (DLMF, 2025, Equations 13.2.39 and 13.7.2). Together, these give the following expansion:
$_{1}F_{1}(a,b;-x)=\frac{\Gamma(b)\,x^{-a}}{\Gamma(b-a)}\,\sum_{k=0}^{\infty}\frac{(a)_{k}\,(1+a-b)_{k}}{k!\,x^{k}},$ (20)
where $(\cdot)_{k}$ is the Pochhammer symbol. In our implementation, we switch to this expansion whenever $|x|\geq 50$ (using the GSL implementation otherwise), and we cut off the sum at five terms. This gives fractional errors of less than $10^{-14}$ for all $x$ of interest when compared against the Boost Math library implementation. [^3] See the Boost library’s excellent discussion of the methods they employ to calculate this function accurately: https://www.boost.org/doc/libs/latest/libs/math/doc/html/math_toolkit/hypergeometric/hypergeometric_1f1.html.
Another important implementation detail is the computation of the overlap integral $g_{ij}$. It is too expensive to compute as needed at runtime, but it can be tabulated ahead of time for the chosen softening kernel, as it depends only on the distance between the two particles and the ratio of their softening lengths. This is done via a technique similar to that described by Fischer et al. (2025); details are given in Appendix D.
Our implementation is open-source under the GNU General Public License and is publicly available on GitHub. [^4] https://github.com/cmhainje/gizmo-public
II.5 Comparison to existing methods
The method we use to treat dark matter bears many similarities to methods for simulating self-interacting dark matter (e.g., Rocha et al., 2013). Our treatment of baryons borrows much technology from analyses of dark matter–baryon interactions in linear cosmology (e.g., Dvorkin et al., 2014; Ali-Haïmoud, 2019). At the time of publication of this article, however, the only existing method specifically designed for simulating dark matter–baryon interactions is Fischer et al. (2025), hereafter F25. We give in this section a detailed comparison of our method to the F25 scheme for “rare, large-angle scatters”.
At a high level, F25 and this work treat dark matter the same way. In both methods, dark matter particles scatter with neighboring baryon particles with probability $(M_{B}/m_{B})\,g_{ij}\,\sigma\,v\,\Delta t$. Further, both sample the scattering baryon velocity $\bm{v}_{B}$ from the Maxwell-Boltzmann distribution described by the bulk velocity and internal energy of the neighbor gas particle.
The key difference is our treatment of baryons. We treat baryons by using the requirement that their velocity distribution remain Maxwell-Boltzmann at all times to calculate moments of the Boltzmann equation, obtaining momentum and heat transfer rates that can be directly applied to the gas macroparticles. F25 instead develop a scheme to apply the probabilistic scatters to the baryons. The issue they must solve, and which we sidestep, is how to conduct a Monte Carlo scattering between two different particle species when the macroparticles represent different numbers of underlying particles: specifically, when the dark matter weight $M_{\chi}/m_{\chi}$ differs from the baryon weight $M_{B}/m_{B}$. They do so by constructing a “virtual” baryon particle with mass $M_{\rm virt}$ set so that $M_{\rm virt}/m_{B}=M_{\chi}/m_{\chi}$, giving both macroparticles equal weight and making scattering straightforward. After scattering, the changes in momentum and energy of the virtual particle are distributed back to the simulation baryon particle, guaranteeing conservation of energy and momentum on a per-scattering basis.
The difficulty with this method is ensuring that the internal energy of the simulation particle remains positive after applying the change in energy of the virtual particle. In particular, to ensure that scatters with all possible sampled baryon velocities $\bm{v}_{B}$ and scattering angles $\Omega$ leave the baryon macroparticle with positive energy, one needs the virtual baryon mass $M_{\rm virt}$ to be only a fraction of the baryon macroparticle mass $M_{B}$. For a given ratio of true particle masses, $m_{\chi}/m_{B}$, this sets an upper limit on the dark matter macroparticle mass $M_{\chi}$ and, when the dark matter mass is small, $m_{\chi}\lesssim m_{B}$, can mean requiring that $M_{\chi}$ be much smaller than $M_{B}$. Such a requirement would be computationally infeasible in many cosmological and galaxy simulations, where the dark matter is the dominant component of the simulation by mass and where high resolution for the baryons is needed. If this mass ratio is not satisfied, F25 note that one can simply reject the fraction of scatters which would lead to negative internal energies. However, they also note that, even though these are a small fraction of all possible scatters, they yield the largest momentum and energy exchange, so cutting them could lead to an underestimate of the effects of dark matter–baryon interactions.
Our method, by divorcing the treatment of the baryons from that of the dark matter, necessarily gives up conservation of energy and momentum on a per-scattering basis, although we prove conservation in the expectation values. However, this split treatment does not impose any restrictions on the masses of the macroparticles, which makes accurate simulation feasible for low mass dark matter.
There is one other, smaller difference which comes about due to our differing treatment of the baryons. The random kicks received by the baryon macroparticles in F25 are applied to the particle velocities $\bm{V}_{j}$, which the particles use as estimates of the local bulk velocity. Since the kicks are randomly oriented, they add variance to the bulk velocity field, effectively siphoning energy out of the internal energies of the particles and into their velocities. In order to prevent a drastic under-reporting of the internal energies of the particles, F25 use viscosity in the baryons to ensure that these random bulk motions are absorbed into the internal energies. To contrast, our approach requires no consideration of viscosity, because the momentum and energy transfer is already split into bulk and thermal components.
III Tests
We have verified our implementation by constructing a test problem that isolates the dark matter–baryon interactions. Explicit details and results are given in Appendix E; we present a brief summary here. The test problem consists of a box with periodic boundary conditions filled uniformly with dark matter and gas of specified densities. The gas particles start at rest and are given some initial internal energy. The dark matter particles are given velocities sampled from a Maxwell-Boltzmann distribution with some bulk velocity and temperature. All physics modules in GIZMO are disabled except for dark matter–baryon interactions. We test a wide variety of initial conditions and mass resolutions for our fiducial dark matter model, which has mass $m_{\chi}=2\,m_{p}$ and velocity-independent cross section $\sigma=10\ {\rm mb}$ ($10^{-26}\ {\rm cm}^{2}$). We also test alternative dark matter models for a limited set of initial conditions, including different dark matter masses, $m_{\chi}\in\{10\,\text{GeV},100\,\text{MeV},10\,\text{MeV}\}$, and velocity-dependent cross sections, $n\in\{-2,-4\}$.
The test problem is designed to facilitate comparison to a direct numerical integration of a system of differential equations derivable under the assumption that the dark matter is Maxwell-Boltzmann distributed at all times. Of course, this assumption does not and should not be expected to hold, as the dark matter departs from equilibrium while transferring energy and momentum to the gas. However, for many choices of the test setup, we find that the dark matter velocity distribution does not stray too far from a thermal distribution, and so its behavior agrees well with the Maxwell-Boltzmann calculation. For other setups where the dark matter evolution departs substantially from a thermal distribution, we cannot directly predict the evolution of the system, but we still verify that the two species equilibrate sensibly and that momentum and energy are conserved. A detailed discussion of our test choices, the comparison to the naïve Maxwell-Boltzmann prediction, and the results of our simulations is given in Appendix E.
Coincidentally, our test setup is almost identical to that which was employed in Fischer et al. (2025). As such, we re-create the most analogous of their tests (which is detailed in their Section 3.4). Our code produces strikingly similar behavior to what they report.
IV Galaxy simulations
Figure 1: Top-down view of the gas in the center of the disk (left) and the entire stellar disk (right) at a variety of cross sections (increasing from top to bottom) and times (increasing from left to right). The largest two cross sections show the development of an extremely compact and dense object in the central gas, as well as the concentration of stars in the center. No significant differences are seen between CDM and the smaller cross sections.
Halo (NFW)
Virial mass $M_{200}$ $1.074\times 10^{12}\ M_{\odot}$
Virial radius $R_{200}$ 205.5kpc205.5\ $\text${kpc}
[…]
Concentration $c$ $10$
Spin parameter $\lambda$ $0.04$
Particles $N_{h}$ $10^{5}$
Stellar bulge (Hernquist)
Mass $M_{b}$ $4.297\times 10^{9}\ M_{\odot}$
Particles $N_{b}$ $1.25\times 10^{4}$
Disk (exponential)
Mass $M_{d}$ $4.297\times 10^{10}\ M_{\odot}$
Scale length $r_{d}$ 3.432kpc3.432\ $\text${kpc}
Scale height $z_{d}$ 0.3432kpc0.3432\ $\text${kpc}
Gas fraction $f_{\text{gas}}$ 0.2
Gas temperature $10^{4}\ \text{K}$
Gas metal fraction $0.02041$
Stellar particles $N_{d,*}$ $10^{5}$
Gas particles $N_{d,\text{gas}}$ $10^{5}$
Table 1: Parameters describing the galaxy initial conditions, from Kim et al. (2016).
As an initial application of our code to a physical system, we simulate an isolated, $z\sim 1$ Milky Way-like galaxy, composed of an NFW profile dark matter halo of virial mass around $10^{12}\ M_{\odot}$, an exponential disk of both stars and gas, and a Hernquist profile stellar bulge. The initial conditions we use are taken from the AGORA isolated disk galaxy test (Kim et al., 2016), where the detailed parameters are presented and explained; we give a summary in Table 1.
With these initial conditions, we perform a suite of simulations using our modified version of GIZMO (Hopkins, 2015) to include dark matter–baryon interactions. We choose a dark matter particle mass of $m_{\chi}=2\,m_{p}$ and velocity-independent cross sections ($n=0$) of $\{0,10,30,100,300,1000\}\ \text{mb}$. We match the parameters and physics used in Kim et al. (2016), including radiative cooling via the Grackle library (Smith et al., 2017), metal production, star formation, a simple thermal stellar feedback model, a Jeans pressure floor (Robertson and Kravtsov, 2008), and gravitational softening lengths of 224 pc for all particle types. For the kernel used, this choice of softening length means gravity is exactly Newtonian at distances greater than 80 pc.
Additionally, we use adaptive softening lengths for the gas. For dark matter, we follow the SIDM implementation in GIZMO (Rocha et al., 2013), which uses adaptive softening lengths only for determining which particles overlap and for the calculation of the overlap integral $g_{ij}$. Importantly, the gravity solver uses fixed softening lengths for dark matter. In preliminary tests, we saw little change when using fixed dark matter softening lengths for the dark matter–baryon interactions. Adaptive softening lengths are initialized to 224 pc, and can vary down to a minimum of 20% of this value (44.8 pc).
We evolve the simulations for approximately 1 Gyr. Since we do not attempt to simulate the cosmic baryon cycle, the gas in the galaxy gets used up for star formation, shutting off the dark matter–baryon interactions prematurely and therefore setting a limit on the length of the simulation. While longer simulations would be interesting, especially for smaller cross sections where the impact of interactions takes longer to build up, these would require a model for gas accretion during the evolution of the galaxy.
IV.1 Impacts on morphology
The primary effect of the dark matter–baryon interactions is the accumulation of both gas and dark matter in the galaxy center when the cross section is large. This increased density can be seen by eye in the gas and stellar disk, which are shown face-on at a few times during the simulations in Figure 1. The gas within the central $\sim 2$ kpc is shown on the left; the entire stellar disk is shown on the right. Each subfigure shows a grid of images, with the cross section increasing from top to bottom and the time increasing from left to right.
The largest cross sections (300, 1000 mb) show that a considerable amount of gas falls into the center, forming an extremely compact and dense central disk of size $\sim 50\ {\rm pc}$. Because this size is approximately the minimum allowed softening length for gas particles and because we use a Jeans pressure floor, higher resolution is required to resolve the structure in the center. For 1000 mb, this process starts sooner and proceeds more quickly, eventually leading to the development of a small ring around the center with a radius of a few hundred parsec. The stellar disk also concentrates in the innermost $\sim 1\text{--}2\ {\rm kpc}$; because stars are collisionless, their response demonstrates that the potential changes drastically. Smaller cross sections show no significant differences from CDM.
Figure 2: Dark matter density profile at the end of the simulations (at about 1 Gyr). Plotted for reference are an NFW profile (dashed grey line) and a profile proportional to $r^{-2}$ (dotted grey line). The largest two cross sections (300 and 1000 mb) are much denser in the center than all other simulations, with 1000 mb becoming steeper even than $r^{-2}$.
The accumulation of dark matter into the center can be seen in the final dark matter density profile, plotted in Figure 2. To facilitate comparison, an NFW profile and a line with slope $r^{-2}$ are shown as well. All profiles are approximately the same for radii outside $\sim 5\ \text{kpc}$ and closely follow the NFW profile. The 300 and 1000 mb simulations show increases in the central density as well as a steepening of the central profile. In particular, the 1000 mb simulation appears to steepen beyond $r^{-2}$ in the center. For cross sections below 300 mb, we see no significant differences from the CDM simulation.
Figure 3: Enclosed dark matter (dashed) and gas (solid) mass profiles at the end (about 1 Gyr) of the simulations. Gas appears to concentrate in the center, with density increasing as a function of cross section. At the largest cross sections (300, 1000 mb), dark matter concentrates in the center as well.
The concentrations of gas and dark matter into the center can also be compared by looking at their enclosed mass profiles, which are shown in Figure 3. All gas profiles are roughly the same outside a radius of about 1 kpc; all dark matter profiles appear the same outside of about 5 kpc. But for the highest cross sections, the gas is concentrated at small radius and surrounded by a cavity of around 1 kpc in size, as seen in the face-on visualizations of Figure 1. This concentration of baryons is mirrored in the dark matter profiles. We investigate possible mechanisms driving these inflows—bar formation and transfer of angular momentum between the gas and dark matter—in Sections IV.3 and IV.4.
IV.2 Impacts on gas and stars
Figure 4: The median gas temperature computed in 1 kpc bins of cylindrical radius. For a cross section of 1000 mb, there is a significant temperature increase at all radii, but the effect is largely gone for cross sections below 300 mb.
Another quantity of interest is the temperature profile of the gas disk. For the final snapshot of each simulation, we compute the median gas temperature in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$; the temperature profile is shown in Figure 4. We see no significant differences at small cross sections. With the 1000 mb simulation, however, the gas temperature is increased by a significant amount at most radii. We note that this is expected: in a Milky Way-like galaxy, the dark matter velocity dispersion is of order a few hundred km/s, which translates to an effective temperature of $\mathcal{O}(10^{6}\,\text{K})$ for $m_{\chi}=2\,m_{p}$. Hence, the dark matter is “hotter” than the gas and therefore heats it for sufficiently strong interactions.
Figure 5: Star formation rate over 1 Gyr for all simulated cross sections. For a cross section of 1000 mb, the star formation rate is increased at nearly all times. At other cross sections, there are few significant differences.
Because dark matter–baryon interactions both heat the gas and change its density, the effects on the star formation history of the galaxy may be non-trivial. To investigate that, for each snapshot we track the total mass in newly-formed stars, and take $\Delta M_{\star}/\Delta t$ as an estimate of the star formation rate. Figure 5 shows the evolution of the star formation rate over the history of the simulation. With 1000 mb, we see enhanced star formation rates starting almost immediately, despite the elevated gas temperature. At 300 mb, the star formation rate shows few differences from CDM until around 600 Myr, at which there is a short burst of increased star formation. At smaller cross sections, there are only small deviations from CDM which are uncorrelated with cross section, and so are likely just stochastic noise.
IV.3 Possible role of bar formation
Figure 6: Bar strength of the stellar disk as a function of time. The bar strength is computed as the maximum of the second Fourier mode (Equation 21) over 1 kpc bins of cylindrical radius spanning $R\in[0,15]\ {\rm kpc}$. A strength of 0.2 is a strong indicator of bar formation.
Because bar formation is known to drive gas inflows toward the galaxy center (Fanali et al., 2015), it is important that we disentangle whether dark matter–baryon interactions or a possible bar is responsible for the increased concentration discussed in Section IV.1. To this end, we investigate whether a bar forms in this simulation. The strength of a bar can be quantified using the second Fourier mode:
$b=\frac{\big|\sum_{j}M_{j}\,\exp(2i\,\varphi_{j})\big|}{\sum_{j}M_{j}},$ (21)
where $M_{j}$ is the mass of star particle $j$ and $\varphi_{j}$ is its azimuthal angle (Beane et al., 2023). For every snapshot of the simulations, we compute the bar strength $b$ in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$ and record the maximum over the bins. Figure 6 shows the bar strength as a function of time, revealing that it evolves almost identically across all cross sections, reaching values above 0.2—a strong indicator for bar formation—within 300 Myr. In particular, during the period of bar formation (the first 300 Myr), we find no dependence of either the bar strength or speed of bar formation on cross section for these initial conditions.
Figure 7: Enclosed dark matter (dashed) and gas (solid) mass profiles at the end of simulations with the stellar disk frozen in order to prevent bar formation. These simulations behave similarly to the simulations with bar formation, indicating that the inflow of gas and dark matter to the galaxy center is not the result of bar formation.
Having convinced ourselves that formation of the bar is not affected by the presence or absence of dark matter–baryon interactions, we next isolate the dark matter–baryon effects from bar-driven inflows. To do this, we perform a set of simulations with the stellar disk frozen in place; since the stars do not move, they cannot form a bar. The enclosed mass profiles of the end states of these simulations are shown in Figure 7. The results are very similar to the simulations without bar formation, indicating that dark matter–baryon interactions, and not bar formation, are driving gas and dark matter toward the center in this simulation. This does not mean that dark matter–baryon interactions have no effect on bar formation, and we believe this is an interesting question which merits future study.
IV.4 Angular momentum
Figure 8: Evolution of the $z$-component of the total angular momentum for particles in the innermost 15 kpc of the galaxy. “Old stars” denote the stellar disk and bulge which are part of the initial conditions; “young stars” denote those which are formed out of the gas during the simulation. At large cross sections, the dark matter gains angular momentum relative to the CDM simulation, while the gas and young stars lose angular momentum. No significant differences can be seen between CDM and the smaller cross sections.
The angular momentum of the gas disk (in particular, its $z$-component) may play a role in the development of gas inflows. Dark matter–baryon interactions introduce a new possible mechanism for transferring angular momentum into or out of the gas disk. To explore this, we investigate the evolution of the $z$-component of the angular momentum during these simulations. Because the dark matter at large radii in general dominates the total angular momentum, we focus on the particles in the innermost 15 kpc. The evolution of the total angular momentum of each component in the innermost 15 kpc is shown in Figure 8. We see that at large cross section (1000 mb) the dark matter gains angular momentum relative to the CDM simulation, while the gas and newly-formed stars lose angular momentum.
Figure 9: Fractional change in the magnitude of the total angular momentum $\bm{L}$ as a function of time. At all cross sections except 1000 mb, we see extremely similar evolution, accumulating about 0.03% additional angular momentum during the simulation, indicating that there is some conservation error as a baseline. At 1000 mb, the accumulation is about twice as large, although it is not clear whether this is due to the implementation or due to the same errors simply exacerbated by the more extreme morphology of the galaxy that develops.
In a galaxy simulation, it is difficult to test for conservation of energy, because physics modules like radiative cooling allow for the dissipation of energy over time. However, angular momentum is conserved, which gives us a useful diagnostic for whether our treatment of the dark matter–baryon interactions has introduced a source of non-conservation. For each snapshot of our simulations, we compute the total angular momentum $\bm{L}=\sum_{i}M_{i}\,\bm{r}_{i}\times\bm{v}_{i}$ and evaluate how the magnitude of this vector evolves. The evolution is shown in Figure 9, revealing that there is a non-zero but small accumulation of angular momentum during the simulation across cross sections: about 0.03% per Gyr, even with dark matter–baryon interactions disabled, or about double that for the 1000 mb case. Some small angular momentum non-conservation may be expected due to the finite separation between macroparticles at the time of interactions (as is true for SIDM implementations, see Tulin and Yu, 2018), which would naturally be greater with large cross sections due to the increased number of scatterings. However, since there is some accumulation in the CDM simulation, it is not clear how much of the increase at 1000 mb should be attributed to our implementation versus the evolution of the strongly reshaped galaxy, which has far more extreme dynamics.
V Summary and Conclusions
We have developed a method for incorporating dark matter–baryon interactions in cosmological $N$-body simulations which is accurate even in the challenging domain of dark matter with comparable or lower mass than the baryon. Our method treats phenomenological power-law cross sections of the form $\sigma=\sigma_{0}\,(v/c)^{n}$. We have implemented the method in the simulation code GIZMO (Hopkins, 2015). Our implementation is open-source, released under the GNU General Public License, and available on GitHub. [^5] https://github.com/cmhainje/gizmo-public We tested the accuracy of the model in a range of environments, including one test which is reproduced from Fischer et al. (2025) and reveals nearly identical behavior.
We have applied this implementation to the evolution of an isolated Milky Way-like disk galaxy, using initial conditions from the AGORA project (Kim et al., 2016) and simulating the system for 1 Gyr. We considered a dark matter particle of mass $2\,m_{p}$ with an isotropic, velocity-independent DM–baryon cross section of 0, 10, 30, 100, 300, and 1000 mb. For the cases of 300 and 1000 mb, the DM–b interaction makes a clear change to the mass distribution in the central kiloparsec of the galaxy, significantly increasing the concentration of baryons and dark matter at the galaxy center in less than 1 Gyr. This occurs both with and without bar formation. Interactions at this strength also heat the gas, increase the star formation rate, and facilitate the transfer of angular momentum out of the gas disk.
In our simulations with cross sections of 100 mb and less, the initial gas supply is largely used up before the DM–b interactions have any significant impact. This is not surprising since we do not simulate gas accretion into the galaxy. Properly simulating the effect of a DM–b cross section of order 10 mb or less—the range relevant for sexaquark dark matter, for instance—demands the use of high-quality hydrodynamics and detailed study of the fidelity of the subgrid models of stellar feedback. These efforts are not justified in this exploratory investigation of an idealized, isolated galaxy. In future work, we plan to study cosmological galaxy formation in the interesting cross section range, with the required sophisticated treatment of the “gastrophysics”.
The code and method we have developed and presented here to incorporate dark matter–baryon interactions gives a practical, efficient, and reliable tool for the community. It should enable robust limits to be placed on various scenarios for dark matter–baryon interactions based on properties of galaxies.
We thank Michael Blanton, Shy Genel, Chris Hayward, David W. Hogg, Clark Miyamoto, and Ben Wandelt for helpful discussions. We especially thank Angus Beane for his advice and debugging aid, and Phil Hopkins for illuminating conversations about GIZMO and the possibilities for implementing this new feature. CH is supported by the National Science Foundation Graduate Research Fellowship under Grant Number DGE-2234660. This research of GRF was supported in part by the Simons Foundation. This work made use of New York University IT High Performance Computing resources, services, and staff expertise.
Appendix A Derivation of dark matter evolution
We begin with the Boltzmann equation for dark matter particle $i$. As written in Equation 8, the single-particle df satisfies
$\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}(\bm{v}_{\chi})=\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,\big[f_{\chi}^{i}(\bm{S}_{\chi})\,\hat{f}_{B}(\bm{S}_{B})-f_{\chi}^{i}(\bm{v}_{\chi})\,\hat{f}_{B}(\bm{v}_{B})\big].$ (A1)
The Boltzmann equation is made of the difference of two rates: the rate of scatters into a state with velocity $\bm{v}_{\chi}$ and the rate of scatters out of velocity $\bm{v}_{\chi}$. The rate of scatters “in” can be conveniently reformulated by instead integrating over all possible initial velocities $\bm{v}_{\chi}^{\prime},\bm{v}_{B}^{\prime}$ and scattering angles $\Omega$, conditioned on the requirement that the post-scatter velocity is $\bm{v}_{\chi}$:
$\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}(\bm{v}_{\chi})=\int{\rm d}^{3}v_{\chi}^{\prime}\,{\rm d}^{3}v_{B}^{\prime}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi^{\prime}B^{\prime}}\,f_{\chi}^{i}(\bm{v}_{\chi}^{\prime})\,\hat{f}_{B}(\bm{v}_{B}^{\prime})\,\delta\big(\bm{S}_{\chi}(\bm{v}_{\chi}^{\prime})-\bm{v}_{\chi}\big)-\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,f_{\chi}^{i}(\bm{v}_{\chi})\,\hat{f}_{B}(\bm{v}_{B}).$ (A2)
We will primarily use this form of the Boltzmann equation in the following derivations.
Consider the evolution of this single-particle df over a timestep $\Delta t$. To first order in the timestep, the update is
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t$ (A3)
The updated df has the form $n_{i}(\bm{x})\,\nu_{i}^{\prime}(\bm{v})$, where $\nu_{i}^{\prime}$ is the new velocity distribution of the particle. Plugging in $\hat{f}_{B}$, we thus find
$n_{i}(\bm{x})\,\nu_{i}^{\prime}(\bm{v})$ $=n_{i}(\bm{x})\,\bigg[\bigg(1-\sum_{j}n_{j}(\bm{x})\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\mathcal{G}(\bm{v}_{B}-\bm{V}_{j},T_{j}/\mu_{j})\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{iB}\,\Delta t\bigg)\,\delta(\bm{v}-\bm{v}_{i})$ (A4)
$\qquad\qquad\qquad+\sum_{j}n_{j}(\bm{x})\int{\rm d}^{3}v_{B}\,{\rm d}\Omega\,\mathcal{G}(\bm{v}_{B}-\bm{V}_{j},T_{j}/\mu_{j})\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{iB}\,\Delta t\,\delta(\bm{v}-\bm{S}_{\chi}(\bm{v}_{i}))\bigg].$
The result contains two integrals over ${\rm d}^{3}v_{B}$ for each gas particle $j$. The second is tricky, as the integration variables $\bm{v}_{B}$ and $\Omega$ are used in the scattering transformation $\bm{S}_{\chi}$ inside the Dirac delta function. As such, it is expedient to approximate the integrals instead by Monte Carlo sampling. The integrals can be performed independently for each particle $j$ using single samples from the baryon Maxwell-Boltzmann distribution, which we denote $\bm{v}^{B}_{ij}$, and the angular component of the differential cross section, $\Omega_{ij}$. The result is
$n_{i}(\bm{x})\,\nu_{i}^{\prime}(\bm{v})=n_{i}(\bm{x})\,\bigg[$ $\bigg(1-\sum_{j}n_{j}(\bm{x})\,\sigma(v_{\rm rel})\,v_{\rm rel}\,\Delta t\bigg)\,\delta(\bm{v}-\bm{v}_{i})$ (A5)
$+\sum_{j}n_{j}(\bm{x})\,\sigma(v_{\rm rel})\,v_{\rm rel}\,\Delta t\,\delta(\bm{v}-\bm{S}_{\chi}(\bm{v}_{i};\bm{v}^{B}_{ij},\Omega_{ij}))\bigg],$
where $v_{\rm rel}\equiv|\bm{v}_{i}-\bm{v}^{B}_{ij}|$, and $\sigma(v)$ is the total cross section, given by $\sigma(v)=\int{\rm d}\Omega\,({\rm d}\sigma/{\rm d}\Omega)\big|_{v}$. The total cross section appears because one can decompose the differential cross section as
$\frac{{\rm d}\sigma}{{\rm d}\Omega}\bigg|_{v,\Omega}=\sigma(v)\,p(\Omega\,|\,v),$ (A6)
where $p(\Omega\,|\,v)$ is the probability density of scattering angles conditioned on the relative velocity. The sample $\Omega_{ij}$ is drawn from $p(\Omega\,|\,v_{\rm rel})$, allowing us to approximate ${\rm d}\sigma/{\rm d}\Omega\approx\sigma(v)\,\delta(\Omega-\Omega_{ij})$, and the rest follows.
In order to isolate $\nu_{i}^{\prime}(\bm{v})$ from all $\bm{x}$ dependence, we can integrate both sides over all space. We find
$\nu_{i}^{\prime}(\bm{v})=\Big(1-\sum_{j}P_{ij}\Big)\,\delta(\bm{v}-\bm{v}_{i})+\sum_{j}P_{ij}\,\delta\big(\bm{v}-\bm{S}_{\chi}(\bm{v}_{i},\bm{v}^{B}_{ij},\Omega_{ij})\big),$ (A7)
where we have introduced
$g_{ij}\equiv\int{\rm d}^{3}x\,W(|\bm{x}-\bm{x}_{i}|,h_{i})\,W(|\bm{x}-\bm{x}_{j}|,h_{j})\quad\text{and}\quad P_{ij}\equiv\frac{M_{j}}{\mu_{j}}\,g_{ij}\,\sigma(v_{\rm rel})\,v_{\rm rel}\,\Delta t.$ (A8)
Appendix B Derivation of baryon evolution
Because the baryons are assumed to thermalize on timescales faster than the timesteps of the simulation, we assume that the baryon velocity distribution is at all times Maxwell-Boltzmann. Thus, the evolution of gas particle $j$ is described by the rates-of-change of the bulk velocity and temperature:
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,\frac{{\rm d}f_{B}^{j}}{{\rm d}t}\bigg|_{\bm{x},\bm{v}_{B}}\,\bm{v}_{B},$ (B1)
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,\frac{{\rm d}f_{B}^{j}}{{\rm d}t}\bigg|_{\bm{x},\bm{v}_{B}}\,(\bm{v}_{B}-\bm{V}_{j})^{2}.$ (B2)
As we proceed, we will use the notation $f_{\chi}=f_{\chi}(\bm{x},\bm{v}_{\chi})$, $f_{\chi}^{\prime}=f_{\chi}(\bm{x},\bm{v}_{\chi}^{\prime})$, $f_{B}=f_{B}^{j}(\bm{x},\bm{v}_{B})$, $f_{B}^{\prime}=f_{B}^{j}(\bm{x},\bm{v}_{B}^{\prime})$.
We start with the acceleration. Inserting the Boltzmann equation gives
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}^{\prime}\,{\rm d}^{3}v_{B}^{\prime}\,{\rm d}\Omega\,f_{\chi}^{\prime}\,f_{B}^{\prime}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B^{\prime}\chi^{\prime}}\,\delta\big(\bm{v}_{B}-\bm{S}(\bm{v}_{B}^{\prime})\big)\,\bm{v}_{B}$
$\qquad-\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\bm{v}_{B}.$ (B3)
The Dirac delta function in the first integral can be integrated immediately. Then, the two terms are identical up to a relabeling of primed and unprimed velocities and the replacement of $\bm{v}_{B}$ by $\bm{S}(\bm{v}_{B})$. Using these facts, we find
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\big[\bm{S}(\bm{v}_{B})-\bm{v}_{B}\big]$ (B4)
$=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}^{2}\,\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,\big[\hat{\bm{v}}_{B\chi}^{\prime}-\hat{\bm{v}}_{B\chi}\big].$ (B5)
For fixed $\bm{v}_{\chi},\bm{v}_{B}$, the integral over $\Omega$ can only result in a vector aligned with $\hat{\bm{v}}_{B\chi}$, so we can project these vectors along this direction. In other words, we can make the following substitution in the integrand: $\hat{\bm{v}}_{B\chi}^{\prime}-\hat{\bm{v}}_{B\chi}\mapsto\hat{\bm{v}}_{B\chi}\,\big[\hat{\bm{v}}_{B\chi}\cdot\hat{\bm{v}}_{B\chi}^{\prime}-1\big]$. We now introduce the momentum transfer cross section:
$\overline{\sigma}(v_{\rm rel})\equiv\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\bigg|_{v_{\rm rel}}\,[1-\cos\theta]=\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\bigg|_{v_{\rm rel}}\,[1-\hat{\bm{v}}_{\rm rel}^{\prime}\cdot\hat{\bm{v}}_{\rm rel}].$ (B6)
Thus,
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=-\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,f_{\chi}\,f_{B}\,\overline{\sigma}(v_{B\chi})\,v_{B\chi}\,\bm{v}_{B\chi}.$ (B7)
Now we insert the Maxwell-Boltzmann form of $f_{B}^{j}$. Introducing $\bm{u}\equiv\bm{v}_{\chi}-\bm{v}_{B}$ and $\bm{w}\equiv\bm{v}_{\chi}-\bm{V}_{j}$, we arrive at
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,n_{j}(\bm{x})\int{\rm d}^{3}v_{\chi}\,f_{\chi}\,\bigg[\underbrace{\int{\rm d}^{3}u\,\mathcal{G}(\bm{w}-\bm{u},T_{j}/\mu_{j})\,\overline{\sigma}(u)\,u\,\frac{\bm{u}\cdot\bm{w}}{w^{2}}}_{\equiv\mathcal{A}(\bm{w};T_{j}/\mu_{j})}\bigg]\,\bm{w},$ (B8)
Finally we can insert the df for the dark matter and integrate both sides over space:
$n_{j}(\bm{x})\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $\approx\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,n_{j}(\bm{x})\sum_{i}n_{i}(\bm{x})\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\,(\bm{v}_{i}-\bm{V}_{j})$ (B9)
$\implies\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}$ $=\sum_{i}\frac{M_{i}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\,(\bm{v}_{i}-\bm{V}_{j}).$ (B10)
The heating rate proceeds similarly. We begin by inserting the Boltzmann equation:
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}^{\prime}\,{\rm d}^{3}v_{B}^{\prime}\,{\rm d}\Omega\,f_{\chi}^{\prime}\,f_{B}^{\prime}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B^{\prime}\chi^{\prime}}\,\delta\big(\bm{v}_{B}-\bm{S}(\bm{v}_{B}^{\prime})\big)\,(\bm{v}_{B}-\bm{V}_{j})^{2}$
$\qquad-\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,(\bm{v}_{B}-\bm{V}_{j})^{2}$ (B11)
$=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\big[(\bm{S}(\bm{v}_{B})-\bm{V}_{j})^{2}-(\bm{v}_{B}-\bm{V}_{j})^{2}\big]$ (B12)
$=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\big[(\bm{S}(\bm{v}_{B})-\bm{v}_{B})^{2}+2\,(\bm{v}_{B}-\bm{V}_{j})\cdot(\bm{S}(\bm{v}_{B})-\bm{v}_{B})\big]$ (B13)
We now split this into two terms. The first contains
$\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,(\bm{S}(\bm{v}_{B})-\bm{v}_{B})^{2}=\bigg(\frac{m_{\chi}\,v_{B\chi}}{m_{\chi}+\mu_{j}}\bigg)^{2}\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,(\hat{\bm{v}}_{B\chi}^{\prime}-\hat{\bm{v}}_{B\chi})^{2}=2\,\bigg(\frac{m_{\chi}\,v_{B\chi}}{m_{\chi}+\mu_{j}}\bigg)^{2}\overline{\sigma}(v_{B\chi}).$ (B14)
The second:
$(\bm{v}_{B}-\bm{V}_{j})\cdot\int{\rm d}\Omega\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,(\bm{S}(\bm{v}_{B})-\bm{v}_{B})$ $=-\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,\overline{\sigma}(v_{B\chi})\,v_{B\chi}\,(\bm{v}_{B}-\bm{V}_{j})\cdot\bm{v}_{B\chi}$ (B15)
$=-\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,\overline{\sigma}(v_{B\chi})\,v_{B\chi}\,\big[v_{B\chi}^{2}\,+(\bm{v}_{\chi}-\bm{V}_{j})\cdot\bm{v}_{B\chi}\big].$ (B16)
Putting these in,
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=2\,\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,f_{\chi}\,f_{B}\,\overline{\sigma}(v_{B\chi})\,\bigg[\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,v_{B\chi}^{3}-v_{B\chi}^{3}-v_{B\chi}\,(\bm{v}_{\chi}-\bm{V}_{j})\cdot\bm{v}_{B\chi}\bigg]$ (B17)
$=2\,\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,f_{\chi}\,f_{B}\,\overline{\sigma}(v_{B\chi})\,\bigg[-\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,v_{B\chi}^{3}-v_{B\chi}\,(\bm{v}_{\chi}-\bm{V}_{j})\cdot\bm{v}_{B\chi}\bigg].$ (B18)
Now, like before, we insert the Maxwell-Boltzmann form of $f_{B}^{j}$, change variables, and find
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=2\,\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,n_{j}(\bm{x})\int{\rm d}^{3}v_{\chi}\,f_{\chi}\,\bigg[-\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{w};T_{j}/\mu_{j})+(\bm{v}_{\chi}-\bm{V}_{j})^{2}\,\mathcal{A}(\bm{w};T_{j}/\mu_{j})\bigg],$ (B19)
where we introduced
$\mathcal{B}(\bm{w};\varsigma^{2})\equiv\int{\rm d}^{3}u\,\mathcal{G}(\bm{w}-\bm{u},\varsigma^{2})\,u^{3}\,\overline{\sigma}(u).$ (B20)
Finally we can insert the form of $f_{\chi}$ and integrate both sides over space:
$n_{j}(\bm{x})\,\frac{3}{2\,\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,n_{j}(\bm{x})\,\sum_{i}n_{i}(\bm{x})\,\bigg[(\bm{v}_{i}-\bm{V}_{j})^{2}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg]$ (B21)
$\implies\frac{3}{2\,\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\sum_{i}\frac{M_{i}}{m_{\chi}+\mu_{j}}\,g_{ij}\bigg[(\bm{v}_{i}-\bm{V}_{j})^{2}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg].$ (B22)
As noted by Ali-Haïmoud (2019), $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions if the momentum transfer cross section has a power law dependence on $v$. Specifically, if $\overline{\sigma}(v_{\rm rel})=\sigma_{0}\,(v_{\rm rel}/c)^{n}$ for integer $n\geq-4$, then
$\mathcal{A}(w;\varsigma^{2})$ $=\frac{\alpha_{n}}{3}\,\frac{\sigma_{0}}{c^{n}}\,\varsigma^{n+1}\,{}_{1}F_{1}\left(-\tfrac{n+1}{2},\tfrac{5}{2},-\tfrac{w^{2}}{2\,\varsigma^{2}}\right),$ (B23)
$\mathcal{B}(w;\varsigma^{2})$ $=\alpha_{n}\,\frac{\sigma_{0}}{c^{n}}\,\varsigma^{n+3}\,{}_{1}F_{1}\left(-\tfrac{n+3}{2},\tfrac{3}{2},-\tfrac{w^{2}}{2\,\varsigma^{2}}\right),$ (B24)
$\alpha_{n}$ $\equiv\frac{2^{(n+5)/2}\,\Gamma(3+\tfrac{n}{2})}{\sqrt{\pi}},$ (B25)
where $_{1}F_{1}$ is the confluent hypergeometric function of the first kind.
Appendix C Equivalence of the two approaches
In this appendix, we show that momentum and energy are conserved in expectation value given the two approaches, so the two approaches are statistically equivalent. For dark matter, we have in the end computed the updated velocity distribution of particle $i$ to be
$\nu_{i}^{\prime}(\bm{v})\bigg|_{t+\Delta t}=\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,\bigg(f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\bigg).$ (C1)
At the end of a timestep, we draw a sample from this distribution $\nu_{i}^{\prime}$, which itself involves approximating ${\rm d}f_{\chi}^{i}/{\rm d}t$ by drawing samples from the distributions of baryon velocities and scattering angles. The average change in velocity, then, is given by integrating over all possible samples (i.e., reverting the sampling approximations):
$\mathbb{E}[\Delta\bm{v}_{i}]$ $=\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\Bigg(f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\Bigg)\,\bm{v}-\bm{v}_{i}$
$=\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\,\bm{v}$ (C2)
Notice that the integral over $({\rm d}f/{\rm d}t)\,\bm{v}$ is exactly how we define the mean acceleration for the baryons. One can follow the derivation of ${\rm d}\bm{V}_{j}/{\rm d}t$ in Appendix B line-by-line with the subscripts $B$ and $\chi$ exchanged to obtain this expectation value; the entire derivation is symmetric. The result is
$\mathbb{E}[\Delta\bm{v}_{i}]$ $=\sum_{j}\frac{M_{j}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\,(\bm{V}_{j}-\bm{v}_{i})\,\Delta t.$ (C3)
The change in energy can be similarly considered.
$\mathbb{E}[\Delta E_{i}]$ $=\frac{M_{i}}{2}\,\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\Bigg(f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\Bigg)\,\bm{v}^{2}-\bm{v}_{i}^{2}$ (C4)
$=\frac{m_{\chi}}{2}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\,\bm{v}^{2}$ (C5)
$=\frac{m_{\chi}}{2}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,\big[\bm{S}(\bm{v})^{2}-\bm{v}^{2}\big]\,\Delta t$ (C6)
$=\frac{m_{\chi}}{2}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{\chi B}\,\big[(\bm{S}(\bm{v})-\bm{v})^{2}+2\,\bm{v}\cdot(\bm{S}(\bm{v})-\bm{v})\big]\,\Delta t$ (C7)
$=\frac{m_{\chi}\,\mu_{j}}{m_{\chi}+\mu_{j}}\int{\rm d}^{3}x\,{\rm d}^{3}v\,{\rm d}^{3}v_{B}\,f_{\chi}\,f_{B}\,\overline{\sigma}\,\bigg[-\frac{m_{\chi}}{m_{\chi}+\mu_{j}}\,v_{\chi B}^{3}-v_{\chi B}\,\bm{v}_{B}\cdot\bm{v}_{\chi B}\bigg]\,\Delta t$ (C8)
$=\frac{m_{\chi}\,\mu_{j}}{m_{\chi}+\mu_{j}}\int{\rm d}^{3}x\,{\rm d}^{3}v\,{\rm d}^{3}v_{B}\,f_{\chi}\,f_{B}\,\overline{\sigma}\,\bigg[\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,v_{\chi B}^{3}-v_{\chi B}\,\bm{v}_{\chi}\cdot\bm{v}_{\chi B}\bigg]\,\Delta t$ (C9)
$=m_{\chi}\,\sum_{j}\frac{M_{j}}{m_{\chi}+\mu_{j}}\,n_{j}(\bm{x})\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,f_{\chi}\,\bigg[\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{w};T_{j}/\mu_{j})-\bm{v}_{\chi}\cdot(\bm{v}_{\chi}-\bm{V}_{j})\,\mathcal{A}(\bm{w};T_{j}/\mu_{j})\bigg]\,\Delta t$ (C10)
$=M_{i}\,\sum_{j}\frac{M_{j}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\bigg[\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\bm{v}_{i}\cdot(\bm{v}_{i}-\bm{V}_{j})\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg]\,\Delta t$ (C11)
Suppose we have exactly one dark matter particle $i$ and one gas particle $j$. Then, the rates-of-change satisfy the following relationships:
$M_{i}\,\mathbb{E}[\Delta\bm{v}_{i}]+M_{j}\,\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}\,\Delta t=0,$ (C12)
$\mathbb{E}[\Delta E_{i}]+\frac{3\,M_{j}}{2\,\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}\,\Delta t+M_{j}\,\bm{V}_{j}\cdot\frac{{\rm d}\bm{V}_{j}}{{\rm d}t}\,\Delta t=0.$ (C13)
These relationships are recognizable as conservation of the expectation values of energy and momentum.
As we scale up to many dark matter and many gas particles, these relationships continue to hold for each individual pair. This is because both the full expectation values we have just computed and the baryonic rates-of-change simply sum over neighbors of the other type. Hence, for each pair $(i,j)$, equal and opposite terms appear in the expectation values for the change to $i$ and in the rates-of-change of $j$.
Appendix D Computation of overlap integral
Let us consider the computation of the overlap factor $g_{ij}$. It is defined
$g_{ij}\equiv\int{\rm d}^{3}x\,W(|\bm{x}-\bm{x}_{i}|;h_{i})\,W(|\bm{x}-\bm{x}_{j}|;h_{j}).$ (D1)
We assume that the gas and dark matter are softened with the same kernel $W$; the transformations and tabulation described below will need to be reworked if the two use different softening kernels.
Define the distance $|\bm{x}_{i}-\bm{x}_{j}|\equiv d$. Let us translate our integration variable to place $\bm{x}_{i}$ at the origin. The integral is invariant to rotation, so we choose to place $\bm{x}_{j}$ at $d\,\hat{\bm{z}}$. Then,
$g_{ij}\equiv\int{\rm d}^{3}x\,W(|\bm{x}|;h_{i})\,W(|\bm{x}-d\,\hat{\bm{z}}|;h_{j}).$ (D2)
We now move to spherical coordinates: $\bm{x}\to(r,\theta,\phi)$, so
$g_{ij}$ $=\int{\rm d}r\,{\rm d}c\,{\rm d}\phi\,r^{2}W(r;h_{i})\,W(\sqrt{r^{2}+d^{2}-2\,d\,r\,c};h_{j})$ (D3)
$=2\pi\,\int_{0}^{h_{i}}{\rm d}r\,r^{2}\,W(r;h_{i})\,\int_{-1}^{+1}{\rm d}c\,W(\sqrt{r^{2}+d^{2}-2\,d\,r\,c};h_{j}),$ (D4)
with $c\equiv\cos\theta$.
Generally, the softening kernel depends only on the ratio of its arguments, up to a normalizing factor in terms of $h$. In other words,
$W(x;h)=\frac{1}{h^{3}}\,w(x/h)$ (D5)
for some $w$. Let $s\equiv r/h_{i}$ and $\delta\equiv d/h_{i}$. Then,
$g_{ij}$ $=2\pi\int{\rm d}r\,{\rm d}c\,r^{2}\,W(r;h_{i})\,W(\sqrt{r^{2}+d^{2}-2\,d\,r\,c};h_{j})$ (D6)
$=\frac{2\pi}{h_{i}^{3}\,h_{j}^{3}}\int{\rm d}r\,{\rm d}c\,r^{2}\,w(r/h_{i})\,w(\sqrt{r^{2}+d^{2}-2\,d\,r\,c}/h_{j})$ (D7)
$=\frac{2\pi}{h_{j}^{3}}\,\int_{0}^{1}{\rm d}s\,s^{2}\,w(s)\,\int_{-1}^{+1}{\rm d}c\,w(\sqrt{s^{2}+\delta^{2}-2\,\delta\,s\,c}/(h_{j}/h_{i})).$ (D8)
For a given softening kernel, the integral which remains can be tabulated in terms of two parameters: $\delta$ and $h_{j}/h_{i}$.
Because $g_{ij}$ is symmetric, we need only tabulate values of the ratio $h_{j}/h_{i}$ between 0 and 1; any ratio $h_{j}/h_{i}>1$ can be handled by swapping $i\leftrightarrow j$.
The range of $\delta$ can also be considered. We assume $W$ is zero beyond the smoothing length, so the overlap is zero if $d>h_{i}+h_{j}$. Thus, the range of $\delta$ which gives non-zero overlap can be found:
$d<h_{i}+h_{j}\implies\delta<1+\frac{h_{j}}{h_{i}}.$ (D9)
Since we choose $i$ and $j$ such that $h_{j}/h_{i}\leq 1$, $\delta$ only needs to be tabulated from 0 to $2$. For small values of $h_{j}/h_{i}$, parts of the table will simply be zero.
We note that this method is almost the same as that described by Fischer et al. (2025), although we use spherical rather than cylindrical coordinates to better match the SIDM implementation in GIZMO (Rocha et al., 2013).
Appendix E Test details
As described in Section III, we test our implementation with a simple but fairly generic test problem. We set up a $(10\ {\rm kpc})^{3}$ box with periodic boundary conditions and fill it uniformly with gas and dark matter of densities $\rho_{B},\rho_{\chi}$. The gas cells are initialized at rest on a regular grid with internal energy set to $(3/2)\,(T_{B}^{0}/m_{B})$, for $T_{B}^{0}$ an initial temperature and $m_{B}$ the baryon particle mass, assumed to be constant and equal to $m_{p}$. The dark matter particles are given uniformly random positions and velocities sampled from a Maxwell-Boltzmann distribution with mean $\bm{V}_{\chi}=V_{\chi}^{0}\,\hat{\bm{x}}$ and temperature $T_{\chi}^{0}$. We use $N_{B}$ gas cells and $N_{\chi}$ dark matter particles. We then simulate the system with all physics disabled except for dark matter–baryon interactions.
We perform this test for many choices of these quantities, detailed in Table 2. The standard setup we choose is $\rho_{B}=\rho_{\chi}=10^{10}\ M_{\odot}\,{\rm kpc}^{-3}$, $T_{B}^{0}=10^{4}\ {\rm K}$, $T_{\chi}^{0}=10^{6}\ {\rm K}$, and $V_{\chi}^{0}=200\ {\rm km}\,{\rm s}^{-1}$, which is similar to the conditions near the center of our isolated disk galaxy. The simulations are then analyzed by computing the total momentum (along $x$), the temperature, and the energy of each component for each snapshot.
Name $m_{\chi}$ $\sigma_{0}$ $n$ $V_{\chi}^{0}$ $T_{\chi}^{0}$ $T_{B}^{0}$ $\rho_{B}$ $\rho_{\chi}$ $N_{B}$ $N_{\chi}$
${\rm cm}^{2}$ ${\rm km}\,{\rm s}^{-1}$ K $M_{\odot}\,{\rm kpc}^{-3}$
Fiducial $2\,m_{p}$ $10^{-26}$ 0 $200$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$V_{\chi}=0$ $2\,m_{p}$ $10^{-26}$ 0 $\bm{0}$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$T_{\chi}=T_{B}=0$ $2\,m_{p}$ $10^{-26}$ 0 $200$ $\bm{10^{1}}$ $\bm{10^{1}}$ $^{10} 1010$$^{10}$ $32^{3}$ $32^{3}$
$\rho_{B}=10\,\rho_{\chi}$ $2\,m_{p}$ $10^{-26}$ 0 $200$ $10^{6}$ $10^{4}$ $\bm{10^{9}}$ $10^{10}$ $32^{3}$ $32^{3}$
$\rho_{\chi}=10\,\rho_{B}$ $2\,m_{p}$ $10^{-26}$ 0 $200$ $10^{6}$ $^{4} 1010$$^{10}$ $\bm{10^{9}}$ $32^{3}$ $32^{3}$
$N_{\chi}=24^{3}$ $2\,m_{p}$ $10^{-26}$ 0 $200$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $\bm{24^{3}}$ $32^{3}$
$N_{\chi}=N_{B}=24^{3}$ $2\,m_{p}$ $10^{-26}$ 0 $200$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $\bm{24^{3}}$ $\bm{24^{3}}$
$m_{\chi}=10\ {\rm GeV}$ $\bm{10\,{\rm GeV}}$ $10^{-26}$ 0 $200$ $\bm{5\times 10^{6}}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$m_{\chi}=100\ {\rm MeV}$ $\bm{100\,{\rm MeV}}$ $\bm{2\times 10^{-27}}$ 0 $200$ $\bm{5\times 10^{4}}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$m_{\chi}=10\ {\rm MeV}$ $\bm{10\,{\rm MeV}}$ $\bm{2\times 10^{-27}}$ 0 $200$ $\bm{5\times 10^{3}}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$\sigma\sim v^{-2}$ $2\,m_{p}$ $\bm{2\times 10^{-33}}$ $\bm{-2}$ $200$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
$\sigma\sim v^{-4}$ $2\,m_{p}$ $\bm{5\times 10^{-40}}$ $\bm{-4}$ $200$ $10^{6}$ $^{4} 1010$$^{10}$ $10^{10}$ $32^{3}$ $32^{3}$
F25 §3.4 $\bm{m_{p}}$ $\bm{1.6726\times 10^{-23}}$ 0 $\bm{0}$ $\bm{72.69}$ $\bm{484.6}$ $\bm{10^{7}}$ $\bm{10^{7}}$ $\bm{10^{5}}$ $\bm{21^{3}}$
Table 2: Parameter choices for tests; those which differ from the “fiducial” test are indicated in bold.
A simple numerical prediction for this problem can be made if the dark matter has a Maxwell-Boltzmann distribution at all times. Under this assumption, each population is described only by its bulk velocity and temperature, and the rates-of-change of these quantities are determined by a simple system of differential equations, resulting from taking the first two moments of the Boltzmann equation. (See Ali-Haïmoud (2019) for details.) In this particular test problem, the fact that both species have zero bulk velocity along $y$ and $z$ allows us to simplify the system to the following:
$\frac{{\rm d}V_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{B}-V_{\chi}),$ (E1)
$\frac{{\rm d}V_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{\chi}-V_{B}),$ (E2)
$\frac{{\rm d}T_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}-T_{\chi}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}}{m_{\chi}}\,V_{\chi B}^{2}\right],$ (E3)
$\frac{{\rm d}T_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}-T_{B}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}}{m_{B}}\,V_{\chi B}^{2}\right],$ (E4)
where $V_{\chi B}=|V_{\chi}-V_{B}|$, $v_{\rm th}^{2}=T_{B}/m_{B}+T_{\chi}/m_{\chi}$, and $\mathcal{A}$ and $\mathcal{B}$ are given in Equations 16 and 17. Given choices for the densities $\rho_{B},\rho_{\chi}$, the dark matter mass $m_{\chi}$, the cross section parameters $\sigma_{0},n$, and the initial velocity distribution parameters $V_{\chi}^{0},T_{B}^{0},T_{\chi}^{0}$, this system of equations can be straightforwardly numerically integrated. We call the result of this integration the “naïve Maxwell-Boltzmann prediction.”
The assumption that the dark matter distribution is at all times Maxwell-Boltzmann does not (and should not) hold for the system in this test problem. However, as long as the distribution of dark matter velocities does not depart too far from Gaussian, we see very similar behavior between the simulation and the naïve Maxwell-Boltzmann prediction. To quantify the departure from Maxwell-Boltzmann of the dark matter velocity distribution, we use the Jarque-Bera statistic (Jarque and Bera, 1980):
${\rm JB}=\frac{1}{6}\,\bigg[S^{2}+\frac{1}{4}\,\big(K-3\big)^{2}\bigg],$ (E5)
where $S$ is the skewness and $K$ is the kurtosis of the velocities. (Note that typically this is multiplied by the number of particles, which we forego in order to compare simulations at different resolutions.) Large values of $\rm JB$} mean large departures from Maxwell-Boltzmann. For the analysis of our simulations, we compute $\rm JB$} for the dark matter velocities independently along $x$, $y$, and $z$ for each snapshot. We then take the maximum value seen at any time along any axis as our guide for whether the Maxwell-Boltzmann assumption is satisfactory for a given test.
Figure 10: Global momentum along $x$ (left), temperature (middle), and energy (right) of the components of our test simulations (solid colored) as a function of time. A black dashed line shows the initial value of the total momentum and energy in order to easily see non-conservation. A green dotted line shows a naïve analytic prediction for the evolution of the system based on an assumption that the dark matter distribution is always Maxwell-Boltzmann. The maximum value of the Jarque-Bera statistic is given (“max JB”) as an indicator of how non-thermal the dark matter distribution becomes during the test; larger values indicate larger deviations from Maxwell-Boltzmann. This figure shows the “fiducial” test, run with ten different random seeds. All tests individually show good agreement both with each other and with the M-B prediction, although some scatter is visible due to the stochastic treatment of scatterings in the dark matter. We note that the scatter is larger in the temperature and energy than the momentum, and does not appear to greatly affect the resulting gas temperatures.
Figure 11: As in Figure 10, except varying the initial bulk velocity (top), temperatures (second row), densities (third and fourth), and resolutions (fifth and sixth).
Figure 12: As in Figure 10, except varying the dark matter mass (first three rows) and velocity-dependence of the cross section (bottom two rows).
The simulation results and naïve Maxwell-Boltzmann predictions are summarized in Figures 10 through 12, which show the evolution of the momenta (left panel), temperatures (middle), and energies (right). Each panel shows the simulation results in blue (dark matter), pink (gas), and grey (total) solid lines, the naïve M-B prediction in a dotted green line, and the initial value of the total momentum or energy as a black dashed line. The maximum value of the Jarque-Bera statistic during the simulation is also reported for each test. Figure 10 shows the “fiducial” simulation, run with ten different random seeds in both the initial conditions generator and the seed inside GIZMO. These help to give an idea of the scatter in the results due to the stochasticity of our method. Figure 11 shows the tests which vary the experimental setup (velocities, temperature, densities, resolutions), all using the same dark matter model: mass $m_{\chi}=2\,m_{p}$, velocity-independent cross section $\sigma=10\ {\rm mb}$. Figure 12 shows the remaining tests, which vary the dark matter model but all use the same experimental setup (note that we change the initial dark matter temperature to keep the velocity dispersion fixed).
In all simulations except the reduced dark matter resolution ($N_{\chi}=24^{3}$), we see excellent conservation of the global momentum and energy, and the lack of perfect conservation in this one simulation is consistent with increased stochasticity because of the reduced resolution. All simulations show sensible behavior, with the dark matter and baryons equilibrating in both momentum and temperature. In all cases where the dark matter remains relatively Maxwell-Boltzmann (empirically, when $\text{max JB}\lesssim 0.01$), we see excellent agreement between the simulation and the M-B prediction. For tests that attain larger JB values, the M-B prediction deviates from the simulation, usually predicting faster energy and momentum transfer than is observed in the simulation. We stress that this is a failure of the naïve Maxwell-Boltzmann prediction and not of the simulation: for these test setups, the dark matter really does deviate from a Maxwell-Boltzmann distribution. Importantly, this suggests that assuming a Maxwell-Boltzmann distribution can be misleading for situations which arise in the evolution of galaxies. This has been similarly pointed out for the effects of dark matter–baryon interactions on linear cosmology observables (Ali-Haïmoud, 2019).
Figure 13: (Top row) Energetics of the “F25 §3.4” test problem, which re-creates the test detailed in section 3.4 of Fischer et al. (2025). These plots are nearly indistinguishable from their Figures 12 and 13 (bottom row).
Serendipitously, Fischer et al. (2025) used an almost identical test setup, and included one test simulation using a constant cross section with isotropic scatterings. It was straightforward to re-create that simulation with our pipeline. The parameters used are listed in Table 2 in the final row, titled “F25 §3.4”: briefly, compared to our fiducial test, the F25 test features proton-mass dark matter, a much larger cross section, smaller densities and temperatures, lower resolution, and a longer simulation time. Figure 13 shows the results. Because the bulk velocities in this problem are zero, the global momentum plot contains no information and the temperature and energy plots are entirely redundant, so we show only the energies. We also include a zoomed-in plot of the dark matter energy and its ratio versus the naïve Maxwell-Boltzmann prediction. The behavior is indistinguishable from that reported in Figures 12 and 13 of Fischer et al. (2025), including the deviation of the simulated result from the naïve Maxwell-Boltzmann-estimated dark matter energy.